### v1

# Boosting Adverse Weather Crowd Counting via Multi-queue Contrastive Learning

Abstract
Currently, most crowd counting methods have outstanding performance under normal weather conditions. However, they often struggle to maintain their performance in extreme and adverse weather conditions due to significant differences in the domain and a lack of adverse weather images for training. To address this issue and enhance the model’s robustness in adverse weather, we propose a two-stage crowd counting method. Specifically, in the first stage, we introduce a multi-queue MoCo contrastive learning strategy to tackle the problem of weather class imbalance. This strategy facilitates the learning of weather-aware representations by the model. In the second stage, we propose to refine the representations under the guidance of contrastive learning, enabling the conversion of the weather-aware representations to the normal weather domain. While significantly improving the robustness, our method only marginally increases the weight of the model. In addition, we also create a new synthetic adverse weather dataset. Extensive experimental results show that our method achieves competitive performance.
Index Terms:
Crowd counting, adverse weather, contrastive learning.
I Introduction
Crowd counting has attracted much attention recently due to its wide range of applications such as public safety, video surveillance, and traffic control. Currently, most of the crowd counting methods [1, 2, 3] are able to estimate the number of crowds well on the images recorded under normal weather conditions. However, when it comes to adverse weather conditions such as rain, haze, and snow, these methods face challenges in maintaining their performance due to domain differences and the limited availability of adverse weather images. For example, in the JHU-Crowd++ dataset [4], the number of images under rain conditions accounts for only 3% of the total dataset.
Figure 1: Trade-off of model weight and accuracy between our proposed method and state-of-the-art methods under adverse weather (a) and normal weather (b) conditions on the JHU-Crowd++ dataset. The radius of the circle is proportional to the number of parameters of the model.
To mitigate the negative influence of adverse weather, an intuitive remedy is to pre-process the images using image restoration modules before counting. Unfortunately, even when restoration modules can mitigate the obscuration caused by adverse weather, the restored images still exhibit significant domain differences from normal weather images. Moreover, the additionally introduced classification and enhancement modules significantly increase the computational burden. Training a separate model for each weather category is also not feasible because the limited samples of adverse weather conditions are insufficient to support the training. [5] introduces a transformer-based approach that addresses the issue of generating complementary information from image-specific degradation without the need for customized image enhancement modules. However, this transformer-based approach not only poses a heavy computational burden but also ignores the issue of weather class imbalance in the dataset.
The aim of this study is to enhance the robustness of crowd counting models in unknown adverse weather conditions while maintaining their performance under normal weather conditions without significantly increasing model complexity. To achieve this objective, we approach the problem of crowd counting, which includes both normal and adverse weather conditions, as an imbalanced multi-domain learning task. The key prerequisite for a model to tackle multi-domain learning is to have perception capabilities across different domains, meaning that the extracted information from each domain should possess discriminative characteristics. This aligns with the objective of contrastive learning [6, 7, 8]. Therefore, in this paper, we propose a two-stage method called Multi-queue Contrastive Learning (MQCL). This approach enables the backbone model to directly extract weather-aware representations, which are further refined by a refiner module. In the first stage, we use contrastive learning to distinguish the characteristics of different weather. However, since the class imbalance mentioned above, the gradient of the loss function of the vanilla contrastive method would be dominated by normal weather images, resulting in poor performance of representations and difficulty in refining and counting. To tackle such an imbalance problem, we design a simple yet effective contrastive learning method called multi-queue MoCo, which replaces the standard single queue in MoCo [7] with multiple queues, providing class-balanced key vectors. In the second stage, all normal weather samples in the queues are treated as positive samples of contrastive learning. Thus the refiner can be guided to convert the representations of adverse weather images to the domain of normal images. Benefiting from the effective representation learning, the refiner and the decoder can be designed to be light to meet our target of not significantly increasing the weight. Compared to the backbone model ConvNeXt [10, 11] used in this paper, our method only introduces 15.3% of the extra FLOPs and 12.7% parameters. Comparison of weight and accuracy of MQCL and state-of-the-art methods is shown in Fig. 1(a/b).
In addition, as currently only one publicly available real-world dataset, JHU-Crowd++ [4], contains adverse weather images and annotations, we synthesized a new dataset called NWPU-Weather based on the NWPU-Crowd dataset [12] with rainy and hazy scenes. This dataset aims to facilitate research on crowd counting in adverse weather conditions. Several representative counting networks are benchmarked to provide an overview of the state-of-the-art performance. Codes and the NWPU-Weather dataset are available at: https://anonymous.4open.science/r/MQCL-B46E/.
The main contributions of our paper are concluded as follows.
  • To boost the robustness of the model under adverse weather conditions while maintaining normal-weather performance, we propose a lightweight two-stage method, achieving significant improvement compared to the baseline.
  • To tackle the problem of class imbalance in contrastive learning, we propose a new method called multi-queue MoCo, achieving better performance than vanilla single-queue MoCo.
  • To realize the conversion of representations from the adverse weather domain to the normal weather domain, we propose a refining method. Contrastive learning is utilized to guide the refinement process, enabling the decoder to focus on a single domain.
  • To provide the crowd counting field with more experimental samples in adverse weather, we synthesize a new adverse weather crowd counting dataset called NWPU-Weather. Extensive experimental results show that our method achieves competitive results.
II Related Work
II-A Crowd Counting under Normal Conditions
Up to now, most single image crowd counting methods can be divided into two categories: density-map-based and localization-based crowd counting. Density-based methods aim to generate a density map, the pixel values of which represent the crowd density at that location. The sum of all pixel values in the density map is the estimated total number of the image. MCNN [1] is a pioneer in employing such a method. Benefiting from the multi-column design, MCNN can handle input images of arbitrary size or resolution. CSRNet [2] employs dilated CNN for the back end to deliver larger reception fields and to avoid pooling operations. More recently, Lin et al. proposed a multifaceted attention network [3] to improve transformer models in local spatial relation encoding. [21] redesigns the multi-scale neural network by introducing a hierarchical mixture of density experts. In addition to model architecture, loss function designing [13, 14] is also a focused area of regression-based crowd counting, which enables the models to effectively learn from ground truth.
On the other hand, the target of detection-based methods is to predict head positions. Early methods obtain the positions by post processing from bounding boxes [31], density maps [37, 38]. However, these methods suffer some inevitable problems such as requirements for complex multi-scale representation and the preclusion of end-to-end training. More recently, Song et al. [15] proposed a model called P2PNet to directly predict a set of point proposals to represent heads in an image, being consistent with the human annotation results. They also proposed a new metric called density Normalized Average Precision (nAP) to provide more comprehensive and more precise performance evaluation. CLTR [33] is a concurrent work with P2PNet, introducing the Transformer architecture to the field of crowd localization and proposing a KMO-based Hungarian matcher to reduce the ambiguous points and generate more reasonable matching results.
II-B Crowd Counting under Adverse Conditions
Existing deep-learning-based methods have achieved unprecedented success with crowd counting, but their performance degraded severely under adverse conditions (e.g., adverse weather) due to the disturbance to the brightness and gradient consistency. However, few research efforts have been made into this problem. Additional class conditioning blocks are utilized by [4] to augment the backbone module, which is trained via cross-entropy error using labels available in the dataset. [5] enables the model to extract weather information according to the degradation via learning adaptive query vectors, but the weight of the model is significantly increased due to the introduction of a Transformer-based module. [36] proposes a single-stage hazy-weather crowd counting method based on direction-aware attention aggregation. However, their method only focuses on the performance in hazy scenes and cannot handle various unknown weather conditions.
II-C Contrastive Learning
Contrastive learning [8, 6] has attracted much attention due to its success in unsupervised representation learning. The target of it is to maximize the similarity of the representations between positive pairs while minimizing that of negative pairs. In recent years, there has been a lot of work to tap the potential of contrastive learning. For example, [7] builds a dynamic dictionary with a queue and a moving-averaged encoder to enable large-scale contrastive learning with dramatically low demand for memory. [9] investigates the contrastive loss and adapted contrastive learning to the field of supervised learning.
Despite the significant success achieved by contrastive learning, Assran et al. [23] revealed that in the formulation of contrastive methods is an overlooked prior to learn features that enable uniform clustering of the data. This prior can hamper performance when pretraining on class-imbalanced data. To solve the class imbalance problem, Kang et al. proposed k-positive contrastive loss (KCL) [39], which draws k instances from the same class to form the positive sample set. Jiang et al. presented a framework called Model-Aware K-center (MAK) [40] to sample open-world unlabeled data to supplement samples for the tail-end classes. Balanced contrastive learning [41] tries to mitigate the unbalanced problem from the aspects of re-weighting the gradient and ensuring batch data balance. Targeted supervised contrastive learning (TSC) [43] improves the uniformity of the feature distribution by generating uniformly distributed cluster centers in advance. Hou et al. [42] clusters each head class into multiple subclasses of similar sizes as the tail classes and enforce representations to capture the two-layer class hierarchy between the original classes and their subclasses.
III Proposed Method
In this work, we aim to improve the robustness of the model under multiple adverse weather conditions and maintain good performance under normal weather. Each sample in the training set consists of three components: the input image, the ground truth points of the human heads and a class label representing the class of the weather. Noting that the weather label is not available to the model in the inferencing phase, which requires the model to be able to deal with unknown corruptions. Images under adverse weather represent only a small part of the dataset. Thus, we formulate our problem as an imbalanced multi-domain learning problem.
Figure 2: The architecture of MQCL. The target of the WRL stage is to learn weather-aware representations via unsupervised contrastive learning. After the WRL stage, the supervised contrastive learning is utilized in the CRR stage to realize the refinement of the representations.
III-A Framework Overview
The architecture of our method is illustrated in Fig. 2. As discussed earlier, we aim to directly enable the crowd counting backbone model to learn weather-aware feature representations and then refine it with a light-weight refiner. Thus, the decoder can focus on a single domain and generate high-quality density map, the sum of which is the estimated number. Since the refiner is premised on stable and consistent representation while it keeps evolving and is not stable during the representation learning stage, we divide the training into two stages to separate these two targets, namely Weather-aware Representation Learning (WRL) stage and Contrastive-learning-guided Representation Refining (CRR) stage, respectively.
In the WRL stage, we use contrastive learning to enable the encoder to learn weather-aware representations. The contrastive learning strategy of the WRL stage is based on our proposed multi-queue MoCo, which can tackle the problem of weather class imbalance. The weights of the encoder and the decoder obtained during the WRL stage will be retained for the CRR stage. In the CRR stage, the encoder is freezed and the main task of this stage is to train a refiner. By treating the normal weather samples as positive examples of contrastive learning, the refiner is guided to pull the adverse weather representations towards normal weather domain. Finally, high-quality density map can be generated to realize precise counting.
III-B Weather-aware Representation Learning
The model architecture of the WRL stage is illustrated in the upper half of Fig. 2. The contrastive learning method [8, 6] is utilized to train the encoders and endow it with the capability to extract weather-aware representations, i.e., images with similar weather conditions correspond to similar representations, whereas those with dissimilar conditions correspond to more distant representations.
Our training structure is the same as MoCo [7], which consists of an encoder $\rm E_{Q}$, a momentum-updated encoder $\rm E_{K}$ and a decoder. The encoders $\rm E_{Q}$ and $\rm E_{K}$ extracts representations $\rm R_{Q}$ and $\rm R_{K}$, respectively, from different augmentations $\rm I_{Q}$ and $\rm I_{K}$ of an image. The representations $\rm R_{Q}$ and $\rm R_{K}$ will be further processed by our proposed multi-queue MoCo strategy, which tackles the problem of weather class imbalance and grant the representations with weather-aware nature, to obtain the contrastive loss. The details of multi-queue MoCo will be elaborated in section III-C. To ensure that the representation simultaneously contains crowd information, the target of the decoder is set as generating a density map under the supervision of the Bayesian loss [13] according to the representation $\rm R_{Q}$. The total loss of the WRL stage is:
$\mathcal{L}_{wrl}=\mathcal{L}_{contra1}+\lambda_{1}\mathcal{L}_{bayesian},$ (1)
where $\mathcal{L}_{contra1}$ is the contrastive loss based on the multi-queue MoCo and $\mathcal{L}_{bayesian}$ is the Bayesian loss.
III-C Multi-queue MoCo
Theoretical reasoning and experimental evidence in [23] suggest that contrastive learning has an overlooked prior-to-learn feature that enables uniform clustering of the data and it can hamper performance when training on class-imbalanced data. In the vanilla contrastive learning strategy, positive and negative samples are entirely obtained through random sampling. This strategy can work perfectly under class-balanced conditions but may struggle if the data is imbalanced due to the mismatch between the actual distribution and the model’s prior.
Figure 3: The architecture of multi-queue MoCo. The projection heads project the representations to 1-D vectors. In the multiple queues, each sub queue is of equal length and corresponds to one weather class.
To tackle the problem, we propose multi-queue MoCo, the architecture of which is illustrated in Fig. 3. Similar to most of the contrastive methods, each image undergoes data augmentation multiple times, and the representations originating from the same image as the anchor are treated as positive samples while those from different images are treated as negative samples. This strategy may lead to a situation where samples with same weather conditions from different images are treated as negatives. This seems to be inconsistent with our motivation. But fortunately, research conducted by [22] indicates that contrastive learning has the tolerance to semantically similar negative samples. In light of this, we conducted extensive experiments and the results indicate that such a strategy outperforms the positive/negative partitioning strategy based on weather labels.
Both representations $\rm R_{Q}$ and $\rm R_{K}$ can be regarded as tensors of $\mathbb{R}^{H\times W\times C_{1}}$, where $H$, $W$ and $C_{1}$ are the height, the width and the number of channels of the representation, respectively. To avoid information loss introduced by the contrastive loss and reduce computational complexity, nonlinear projection heads [6] are introduced after encoders to project the representations to 1-D vectors. The projection head first pools the representations to vectors of $\mathbb{R}^{C_{1}}$ and then project them to vectors Q or K of $\mathbb{R}^{C_{2}}$ by introducing a multi-layer perceptron. $C_{2}$ is the dimension of the vectors. We refer to the projection heads following encoder $\rm E_{Q}$ and $\rm E_{K}$ as projection head $\rm P_{Q}$ and $\rm P_{K}$, respectively. They have the same structure but do not share parameters with each other.
In contrast to MoCo, to achieve a uniform distribution of classes within vectors in the memory, we improve the original single queue to a multi-queue structure. The number of sub queues in the multiple queues is equal to the number of classes, with each sub queue having an equal length and exclusively storing vectors Q that match its corresponding class. The multi-queue structure can be considered as a tensor of $\mathbb{R}^{B\times L\times C_{2}}$, where $B$ is the number of classes and $L$ is the length of each sub queue. Immediately when the computation of vector K is completed, it will be pushed into the corresponding sub queue according to its weather class label. With this design, the number of samples of each weather class in the memory becomes equal, aligning perfectly with the uniform prior of the contrastive loss. Moreover, due to the limited number of images from adverse weather in the dataset, multiple samples from the same image may coexist within a sub queue. To avoid treating the above samples negative to each other, we propose to assign a unique index value to each image in the dataset and treat the samples corresponding to the same index as positive. The loss function of our multi-queue MoCo can be calculated as follows:
$\mathcal{L}_{contra1}=\sum_{i\in I}\frac{-1}{|P(i)|}\sum_{p\in P(i)}\log\frac{\exp({\rm Q_{i}}\cdot{\rm K_{p}}/\tau)}{\sum_{a\in\mathbf{A}}\exp({\rm Q_{i}}\cdot{\rm K_{a}}/\tau)},$ (2)
where $I$ is the batch size, $\rm Q_{i}$ is the anchor vector, $P(i)$ is the set of indices of the vectors K originating from the same image with $\rm Q_{i}$, $\mathbf{A}$ is the set of the indices of all of the vectors in the multiple queues and $\tau$ is the temperature.
III-D Contrastive-learning-guided Representation Refining
After the WRL stage, we can assume that the encoder is “mature” enough to effectively extract the weather and crowd information from images. The task of the CRR stage is to train a refiner which can convert the weather-aware representations to the normal weather domain, generating representations of the same size of $\mathbb{R}^{H\times W\times C_{1}}$ and enabling the decoder to focus on a single domain.
The multiple queues obtained during the WRL stage have stored a sufficient number of vectors from various weather conditions. In this paper, we propose to preserve and freeze these multiple queues and employ contrastive learning to guide the refiner to convert the representation. In this stage, all the vectors in the normal-weather sub queue are treated as positive samples and those in other sub queues are treated as negatives. Since there is no longer a need to generate vectors K, the encoder $\rm E_{K}$ and projection head $\rm P_{K}$ are discarded. In order to maintain stable representations, the parameters of encoder $\rm E_{Q}$ are fixed at this stage. The projection head $\rm P_{Q}$ is preserved, fixed, and moved behind the refiner. We denote the vector projected from the refined representation as $\rm Q^{\prime}$. The model structure of the CRR stage is shown in the lower half of Fig. 2. The loss function of contrastive learning in this stage is calculated as follows:
$\mathcal{L}_{contra2}=\sum_{i\in I}\frac{-1}{|\mathbf{N}|}\sum_{p\in\mathbf{N}}\log\frac{\exp(Q^{\prime}_{i}\cdot{\rm K_{p}}/\tau)}{\sum_{a\in\mathbf{A}}\exp(R({\rm Q_{i}})\cdot{\rm K_{a}}/\tau)},$ (3)
where $\rm N$ is the set of the indices of the vectors K in the normal-weather sub queue. The main difference between the $\mathcal{L}_{contra1}$ and $\mathcal{L}_{contra2}$ is that $\mathcal{L}_{contra1}$ treats samples originate from the same image as positive examples, while $\mathcal{L}_{contra2}$ treats all the normal weather samples in the retained multiple queues as positive examples. Similar to the WRL stage, the CRR stage continues to utilize the Bayesian loss to supervise the density map. The overall loss function $\mathcal{L}_{CRR}$ for this stage is calculated as follows:
$\mathcal{L}_{CRR}=\mathcal{L}_{contra2}+\lambda_{2}\mathcal{L}_{bayesian},$ (4)
where $\mathcal{L}_{contra2}$ is the contrastive loss and $\mathcal{L}_{bayesian}$ is the Bayesian loss.
IV Experiments and Discussions
IV-A Experiment Setup
The ConvNeXt [10] is employed as our backbone model and we use the ConvNeXt-T version for simplicity. The structure of ConvNeXt-T is: C = (96, 192, 384, 768), B = (3, 3, 9, 3), where C stands for the number of channels and B is the number of ConvNeXt blocks in each stage. Following [11], we replace the linear layer at the end of the ConvNeXt by an upsampling block to keep the downsampling rate at 8. The encoder $\rm E_{Q}$ and encoder $\rm E_{K}$ both have the same structure as the first three stages of ConvNeXt and the decoder corresponds to the final stage and the upsampling module. The pre-trained weights of ImageNet-22K [24] are loaded as the initial parameters. Two-layer MLPs are employed in the projection heads, the output dimension of which is 2048 and 128. The refiner consists of three ConvNeXt blocks with input and output dimensions of 768, initialized with random parameters.
As mentioned above, each image in a batch undergoes augmentation twice. Specifically, we random crop the image with a size of 256 $\times$ 256, and horizontal flipping is applied for a probability of 50%. In the multi-queue structure, the number of the sub queues is equal to the number of weather classes, and the length of each sub queue is set to 1024. The AdamW optimizer [25] is adopted both in the WRL and CRR stage, the learning rate is scheduled by a cosine annealing strategy and the initial learning rate is $10^{-4}$. The weight decay is set to $10^{-3}$ and the batch size is $16$. $\lambda_{1}$ and $\lambda_{2}$ in the loss function are both set to 10 and the temperature $\tau$ is set to 0.05.
Following the convention of existing works [2, 3], we adopt Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) as the metrics to evaluate the methods.
TABLE I: Quantitative results comparing with the state-of-the-art methods on the NWPU-Weather dataset. The numbers in parentheses represent the rankings of our method.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
CSRNet [2] (CVPR 18) 74.8 200.7 175.3 730.2
BL [13] (ICCV 19) 69.9 248.1 137.1 339.4
DM-Count [28] (NeurIPS 20) 80.6 319.1 153.6 338.2
KDMG [29] (PAMI 20) 108.0 318.6 151.9 328.4
MAN [3] (CVPR 22) 64.1 259.1 105.9 264.1
ConvNeXt [10] (CVPR 22) 69.3 264.0 108.2 286.3
MQCL (Ours) 61.6 (1) 245.5 (2) 93.3 (1) 250.1 (1)
Figure 4: Examples of normal (a), hazy (b) and rainy (c) weather images in the NWPU-Weather dataset.
IV-B NWPU-Weather Dataset
Considering the current scarcity of crowd counting datasets containing adverse weather scenarios and labels, we synthesize an adverse weather crowd counting dataset containing hazy and rainy scenes based on the NWPU-Crowd dataset [12], namely NWPU-Weather. Since the test set of the NWPU-Crowd dataset is not publicly available, we extract part of the original training set as the test set. The first 1-2609 images from the original dataset are used as the training set, the images from 2610 to 3109 are used as the test set, and the validation set still consists of images from 3110 to 3609.
Consistent with the JHU-Crowd++ dataset, our NWPU-Weather dataset keeps the imbalance of weather types, the hazy and rainy scenes are synthesized by a probability of 5%, respectively. The distribution of the weather conditions in the dataset is shown in Table II. We follow the approach outlined in [26] to synthesize the weather scenarios. The intensity, density, and angle of the rain are set as random values. The depth maps required during the haze synthesis process are estimated by ZoeDepth [27] and the intensity of the haze is also random. Examples of the images in the NWPU-Weather images are shown in the Fig. 4.
TABLE II: The distribution of the number of images under different weather conditions in the NWPU-Weather dataset.
Stage Normal Haze Rain
Train 2365 120 124
Val 446 31 23
Test 442 28 30
Total 3253 179 177
Except for the MQCL, we also benchmark several representative counting networks, providing an overview of the state-of-the-art performance, including CSRNet [2], DM-Count [28], KDMG [29], ConvNeXt [10] and MAN [3]. Note that since the model structure of AWCC-Net [5] is not fully elucidated and the code is not available at present, we do not provide the performance. Following the convention of existing works [2, 3], we adopt Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) as the metrics to evaluate the methods. The quantitative results of counting accuracy are listed in Table I. Compared with the baseline ConvNeXt, MQCL exhibits a significant performance improvement under adverse weather conditions. MAE and RMSE are decreased by 13.8% and 12.6%, suggesting that the proposed representation learning and refining strategy do indeed boost the robustness of the model under adverse weather conditions. Moreover, MQCL can also improve the performance under normal conditions, decreasing MAE and RMSE by 11.1% and 7.0%, respectively. We believe that this can be attributed to the contrastive learning strategy employed in this paper, which considers augmentations from the same image as positive examples. This strategy not only assists the model in weather perception but also strengthens the model’s ability to recognize different scenes. Sampling from the same image ensures that the positive examples not only share the same weather conditions but also possess similar scene characteristics. Compared with the previously best method MAN, MQCL has also achieved significant performance improvement. The MAE/RMSE under normal and adverse weather conditions have reduced by 3.9%/5.2% and 11.9%/5.3%, respectively.
TABLE III: Quantitative results comparing with the state-of-the-art methods on the JHU-Crowd++ dataset. The numbers in parentheses represent the rankings of our method.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
SFCN [30] (CVPR 19) 71.4 225.3 122.8 606.3
BL [13] (ICCV 19) 66.2 200.6 140.1 675.7
LSCCNN [31] (PAMI 20) 103.8 399.2 178.0 744.3
CG-DRCN-V [4] (PAMI 20) 74.7 253.4 138.6 654.0
CG-DRCN-R [4] (PAMI 20) 64.4 205.9 120.0 580.8
UOT [32] (AAAI 21) 53.1 148.2 114.9 610.7
GL [14] (CVPR 21) 54.2 159.8 115.9 602.1
CLTR [33] (ECCV 22) 52.7 148.1 109.5 568.5
MAN [3] (CVPR 22) 46.5 137.9 105.3 478.4
AWCC-Net [5] (ICCV 23) 47.6 153.9 87.3 430.1
ConvNeXt [10] (CVPR 22) 52.7 154.9 105.1 561.4
MQCL (Ours) 50.5 (3) 152.0 (4) 96.5 (2) 522.8 (3)
TABLE IV: Comparison of computational complexity and the number of parameters. The computational complexity is measured by FLOPs when inferencing images with the size of 384 $\times$ 384.
Method FLOPs #param
CLTR [33] (ECCV 22) 37.0G 43M
MAN [3] (CVPR 22) 58.2G 31M
AWCC-Net [5] (ICCV 23) 58.0G+ 30M+
ConvNeXt [10] (CVPR 22) 27.0G 29M
MQCL (Ours) 31.2G 32M
Figure 5: The pie chart of the weather condition distribution of the JHU-Crowd++ dataset (a) and the GCC dataset (b).
IV-C JHU-Crowd++ Dataset
There are 4372 images and 1.51 million labels contained in the JHU-Crowd++ dataset [4]. Out of these, 2272 images were used for training, 500 images for validation, and the remaining 1600 images for testing. The advantage of JHU-Crowd++ is its inclusion of diverse scenes and environmental conditions, such as rain, snow and haze. It also provides weather condition labels for each image. Due to the rarity of adverse weather, the weather classes in JHU-Crowd++ are imbalanced. As is shown in Fig. 5 (a), the number of images under rain, snow, and haze conditions accounts for only 3%, 5%, and 4% of the total dataset, respectively.
As shown in Table III, despite the challenges posed by the diverse scenes, complex and variable weather conditions and weather class imbalance in the JHU-Crowd++ dataset, MQCL achieves an improvement of 8.6% in MAE and 6.9% in RMSE under adverse weather conditions compared to the baseline ConvNeXt. This indicates that MQCL not only performs well on synthetic datasets but also effectively boosts the model’s robustness in real-world datasets. However, while MQCL achieved significant improvements over the baseline model ConvNeXt, its performance still slightly lags behind the current state-of-the-art algorithms. We summarize the reasons as follows: 1) as shown in Table III, all methods outperforming MQCL, including CLTR, MAN, and AWCC-Net, have introduced Transformer-based modules, the computational complexity of which is significantly higher than MQCL. The MAE of AWCC-Net under adverse weather conditions is 9.5% lower than our method, but it incurs over 85.9% more FLOPs; 2) MQCL focuses on the design of learning strategies. It only introduces a lightweight refiner module without significantly altering the architecture of the backbone network. The visualizations of ConvNeXt and our MQCL are compared in Fig. 6.
TABLE V: Quantitative results comparing with the state-of-the-art methods on the GCC dataset. The numbers in parentheses represent the rankings of our method.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
CSRNet [2] (CVPR 18) 64.8 180.5 87.4 220.8
BL [13] (ICCV 19) 104.5 252.6 137.2 304.0
DM-Count [28] (NeurIPS 20) 78.4 207.6 103.5 248.5
KDMG [29] (PAMI 20) 94.6 225.0 124.0 280.5
MAN [3] (CVPR 22) 70.1 207.8 94.5 246.4
ConvNeXt-S [10] (CVPR 22) 65.3 184.5 85.5 221.7
MQCL (Ours) 59.7 (1) 175.5 (1) 79.7 (1) 210.5 (1)
Figure 6: Visualizations of our baseline ConvNeXt and our MQCL. The first row are the input images. The second and the third rows are the density maps predicted by ConvNeXt and our method, respectively. The images are sampled from the JHU dataset, including weather conditions of normal, haze, rain and snow.
IV-D GCC Dataset
GTA5 Crowd Counting (GCC) dataset [30] is a fully synthetic large scale dataset, the scenarios of which is derived from the computer game Grand Theft Auto V (GTA5). The dataset was initially proposed for pretraining the model and boosting the performance on the real-world datasets by finetuning or adaptation. Here, we use it as a benchmark to validate the performance of crowd counting models under adverse weather conditions on large-scale data.
There are 15212 images and 7.63 million labels contained in the GCC dataset. Since each location in the dataset encompasses multiple perspectives, [30] provides three strategies for splitting the training and the testing set: random, cross-camera and cross-location. Here we choose the cross-location strategy, which means 75%/25% locations are utilized for training/testing, since this scheme is most effective in validating the robustness of the model against unknown scenes. The weather conditions in the GCC dataset are diverse, comprising a total of seven types: normal, cloudy, rainy, foggy, thunder, overcast and extra sunny. They are evidently unevenly distributed either and the pie chart of the weather distribution is shown in Fig. 5 (b). In our implementation, we randomly selected 10 locations out of the 75 in the training set as a validation set. And it is worth noting that we choose ConvNeXt-S rather than ConvNeXt-T as our backbone model due to the large-scale nature of the GCC dataset.
The quantitative results of our method and several state-of-the-art methods are compared on the GCC dataset, as shown in Table V. As can be seen, our MQCL achieves optimal performance under both normal and adverse weather conditions on all the metrics. Comparing to our baseline ConvNeXt, MQCL decreases the MAE/RMSE under normal and adverse weather conditions by 8.6%/4.9% and 6.8%/5.1%, respectively. This indicates that our MQCL can also outperforms other methods when trained on relatively large-scale datasets and more importantly, our multi-queue MoCo strategy performs well even in the case of higher number of sub queues (7 for example).
IV-E Key Issues and Discussions
The significance of contrastive learning. As discussed earlier, the prerequisite for addressing the multi-domain learning problem is that the model can perceive different domains. If the encoder fails to extract distinctive weather information, the refiner will lack the target for conversion and struggle to fulfill its intended purpose. As shown in Table VI, if contrastive learning is not employed in the first stage to enable the model to learn weather information with discriminative characteristics, and the refiner module is directly introduced afterwards, the performance of the model actually decreases significantly under both normal and adverse weather conditions. We believe this is because the refiner module, trained from scratch without clear targets, may disrupt the crowd information extracted by the original encoder.
TABLE VI: The significance of contrastive learning in our method. “Refine w/o CL” adopts a similar approach to MQCL, except that $\mathcal{L}_{contra}$ and $\mathcal{L}_{supcon}$ are removed from the loss functions, respectively.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
Encoder and decoder only 52.7 154.9 105.1 561.4
Refine w/o CL 56.2 163.1 107.1 585.6
MQCL (Ours) 50.5 152.0 96.5 522.8
TABLE VII: Performance comparison of the two optional strategies on the JHU-Crowd++ dataset. Strategy 1 treats samples with the same weather label as positive examples and strategy 2 treats samples originating from the same image as positive examples.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
Strategy 1 51.4 153.0 99.7 540.4
Strategy 2 (Ours) 50.5 152.0 96.5 522.8
Strategy of positive/negative selection. As mentioned above, there are two optional strategies in the WRL stage: 1) treating samples with the same weather label as positive examples; 2) treating samples originating from the same image as positive examples; We conduct experiments on the JHU-Crowd++ dataset to compare these two strategies. The model performance after the WRL stage using these two strategies is shown in Table VII. Strategy 2 outperforms Strategy 1 in both normal and adverse conditions. We attribute this to the following reasons: 1) even if two samples share the same weather label, their weather conditions may still vary significantly. The practice of minimizing all the representations with the same label is not in line with the target of us; 2) samples from the same image not only share the same weather conditions but also the same scene. Strategy 2 has the potential to enhance the model’s scene recognition capabilities.
Figure 7: The t-SNE [35] visualization of the vectors Q after the WRL stage on the JHU-Crowd++ dataset using memory bank (a), single-queue MoCo (b) and multi-queue MoCo (c), respectively.
TABLE VIII: Performance comparison of the three storage strategies for vectors K on the JHU-Crowd++ dataset.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
Memory Bank 53.7 157.6 104.9 531.1
Single-queue MoCo 54.6 166.5 107.9 589.7
Multi-queue MoCo (Ours) 50.5 152.0 96.5 522.8
The storage strategy for vectors K. There are three optional storage strategies for vectors K: 1) memory bank [34]; 2) single-queue MoCo [7]; 3) multi-queue MoCo (ours). Extensive experiments are conducted on the JHU-Crowd++ dataset to compare the above strategies. The t-SNE [35] visualization of the vectors Q after the WRL stage on the JHU-Crowd++ dataset is shown in Fig. 7. The memory bank strategy suffers from a severe lack of discriminative capacity. The representations learned by the single-queue strategy are also not discriminative enough, especially near the rain weather representations in Fig. 7(b). The multi-queue strategy does not suffer from the aforementioned issues. The significantly dispersed clusters shown Fig. 7(c) illustrate that the properties learned from the multi-queue MoCo represent weather information well. The performance comparison shown in Table VIII also demonstrates that the proposed multi-queue MoCo can effectively address the class imbalance problem. We attribute the phenomenon to the following reasons: 1) while the memory bank can store a large number of samples with minimal memory consumption, it does not employ a stable strategy to update the encoder, and the sample update frequency is too low, resulting in poor sample consistency; 2) although single-queue MoCo addresses the issue of poor sample consistency by introducing a queue and momentum update strategy, the class imbalance problem in the dataset leads to inconsistencies between the data distribution in the queue and the uniform distribution prior; 3) multi-queue MoCo not only retains the advantages of high sample consistency but also greatly alleviates the problem of inconsistency between data distribution and the uniform prior. Thus, it achieves the best performance.
TABLE IX: Ablation study on the JHU-Crowd++ dataset.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
Encoder and decoder only 52.7 154.9 105.1 561.4
+ Multi-queue MoCo 51.2 152.6 97.2 524.6
+ Refiner 50.5 152.0 96.5 522.8
Ablation study. Ablation studies are performed on the JHU-Crowd++ dataset and the quantitative results are shown in Table IX. We start with the baseline of the end-to-end model, i.e., only the encoder and decoder. First, the effectiveness of multi-queue MoCo is tested. An improvement of 2.8%/1.5% and 7.5%/6.6% in MAE/RMSE under normal and adverse weather is achieved compared to the baseline. From this, it can be seen that most of the performance improvements under adverse weather conditions come from the representation learning in the WRL stage. This verifies the effectiveness of the proposed multi-queue MoCo for enhancing robustness under adverse weather conditions. Additionally, the performance improvement under normal weather conditions corroborates the earlier analysis that the strategy that treats different augmentations from the same image as positive samples can aid in strengthening the scene recognition capabilities of the model. Subsequently, the refiner is added, and the best performance is achieved, with a reduction of 1.4%/0.4% in MAE/RMSE under normal weather conditions and 0.7%/0.3% under adverse weather conditions, respectively. This demonstrates that, under the guidance of supervised contrastive learning, the refiner is capable of converting adverse weather representations to the normal domain, enabling the decoder to focus on a single domain, resulting in performance improvements under both normal and adverse weather conditions.
TABLE X: Comparison between multi-queue MoCo and other class imbalance contrastive learning strategies on the JHU-Crowd++ dataset.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
KCL [39] (ICLR 20) 51.1 154.3 97.9 564.6
BCL [41] (CVPR 22) 50.5 154.3 97.2 524.2
TSC [43] (CVPR 22) 50.8 152.1 96.4 522.4
SBCL [42] (ICCV 23) 54.0 159.2 101.1 563.0
Multi-queue MoCo (Ours) 50.5 152.0 96.5 522.8
Comparison with other class imbalance strategies. In this section, we will discuss the strengths and weaknesses of multi-queue MoCo compared to the five representative imbalanced contrastive learning strategies mentioned in Section II-C, and investigate their applicability to our MQCL method. Attributed to MAK’s [40] strategy of sampling from the real world, the diversity of tail-end classes is significantly increased, the issue of class imbalance can be effectively addressed. However, it is not applicable to our crowd counting task because only unlabeled samples can be supplemented, but in crowd counting, head position labels are required for learning. The rest four methods can be applied to our model, we conduct experiments on the JHU-Crowd++ dataset to compare and analyze their performance results.
In the experiment setup, only $\mathcal{L}_{contra1}$ in the MQCL is replaced with the loss functions of the four imbalanced contrastive learning strategies, and all other settings remains unchanged. The results are shown in Table X. The performance of KCL [39] is worse than multi-queue MoCo because KCL samples a sufficient number of positive samples, but the negative samples remain imbalanced. SBCL [42] fundamentally alleviates the issue of class imbalance through the method of subclass partitioning. The experiments in [42] show that SBCL outperforms KCL [39] and TSC [43] in image classification tasks. However, it achieved the lowest performance in this experiment. We think the approach of dividing the normal weather class into multiple subclasses may bring confusion to the CRR stage. Because in the CRR stage, the representations are pulled towards the normal weather domain, which, however, has multiple clusters. BCL [41] combines gradient balancing and batch data balancing; TSC [43] introduces cluster center priors. They achieved performance almost as excellent as multi-queue MoCo. Although these three methods exhibit similar performance, they address the imbalance problem from different perspectives, each with its own significance.
TABLE XI: Generalization ability of ConvNeXt-T trained on the NWPU-Normal dataset and the NWPU-Weather dataset on the test set of the JHU-Crowd++ dataset.
Method Normal Adverse
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
NWPU-Normal to JHU-Crowd 65.9 184.4 144.7 705.2
NWPU-Weather to JHU-Crowd 63.9 163.6 142.8 497.2
The significance of the NWPU-Weather dataset. As described in Section IV-B, we synthesize the NWPU-Weather dataset to validate the models’ performance under adverse weather conditions. However, beyond this, in terms of the significance of proposing this dataset, there is a question: how is the generalization ability of the model trained on the synthetic NWPU-Weather dataset? If the model trained on the NWPU-Weather dataset can’t be generalized to the real world, the contribution of the synthetic dataset is limited. Thus we conduct experiments to validate this question.
In the experiment setup, we prepare two datasets as source-domain datasets. One is our proposed NWPU-Weather dataset. And another, for comparison, we replace all the adverse-weather images in the NWPU-Weather dataset with its corresponding normal-weather raw images. Here we call it NWPU-Normal. To test the generalization ability of the models, the target domain is set as the JHU-Crowd dataset since it is fully real-world. We train ConvNeXt-T on the training set of the two source-domain datasets respectively and directly test their performance on the testing set of the target domain dataset without any finetuning.
As is shown in Fig. XI, compared to the model trained on the NWPU-Normal dataset, that trained on our synthesised NWPU-Weather dataset is 3.0%/11.3% and 1.3%/29.5% stronger in MAE/RMSE under normal and adverse weather, respectively. This validates the significance of our NWPU-Weather dataset: more diversified weather conditions in the training data imparts better robustness to the model. Especially notable is the improvement in RMSE under adverse weather conditions, indicating that the synthesized data particularly excels in enhancing the stability of the model.
V Conclusion and Limitations
In this paper, we propose a contrastive learning-based method called MQCL to tackle the problem of class-imbalanced adverse weather crowd counting and synthesize a new adverse weather crowd counting dataset. To address the dual challenges of image degradation and class imbalance, the multi-queue MoCo is employed to enable the model to learn weather-aware representations. Furthermore, a contrastive-learning-based refining method is utilized to guide the refiner on representation conversion. Extensive experiments are conducted to compare and choose the strategy of contrastive learning. We believe that this is not only applicable to the crowd counting task but also holds significant reference value for other domains.
MQCL has achieved significant performance improvements compared to the baseline. However, there are still some limitations. For example, we focus on the learning strategy and maintaining the model’s lightweight nature, but its performance still falls slightly behind the state-of-the-art models based on Transformers. Secondly, for this study, our focus is on the application of multi-queue MoCo in the adverse weather crowd counting task, but it is a generic method to deal with data imbalance, its performance in other tasks awaits future research. Additionally, we only utilized synthetic methods to construct the dataset, leading to a certain domain gap between real-world scenes.